OpenStudy (anonymous):

In a certain function, y varies directly with x. If y = 6 when x = 8, find x when y = 9.

OpenStudy (anonymous):

can we say y=tx; t being any constant then 6/8=t then when y=9 then x=12 ?? is it ryt?

OpenStudy (anonymous):

12 is one of the options so i guess. thanks alot. there is another one please, can you help. The population of a southwestern town is expected to grow by an average of 711 new residents each year. If the town’s population in the year 2010 was 19,584, write an equation predicting P(t), the town’s population t years after 2010. the options are: P(t) =2010+711 P(t)=19.584(2010+t)+711 P(t)=711+19,584 p(t)=2010t+19,584

OpenStudy (anonymous):

ok thanks bibby

OpenStudy (ikram002p):

cuz y varies directly with xu can do this..